OpenStudy (jaredstone4):

Evaluate: Integral [1,e^2] [ln(x^2) / x] dx

OpenStudy (jaredstone4):

\[\int\limits_{1}^{e^2} [\ln(x^2) / x ] dx\]

OpenStudy (mathmale):

What have you tried so far? ln (x^2) can be written in another form, using one of the "rules of logs": ln a^b = b*ln a. Care to try that?

OpenStudy (jaredstone4):

I really didn't try much to be quite honest. I didn't recognize it as any integral rule right off the bat so I just moved on for the time being. Does it have to become a u-substitution?

OpenStudy (mathmale):

A substitution would definitely make this problem easier. i do suggest that you rewrite ln (x^2) using the appropriate rule of logs. One of the other rules of logs is ln a + ln b = ln (a*b). Hope that triggers your memory of the third rule.

OpenStudy (jaredstone4):

I will. Are you referring to lna - lnb = ln a/b?

OpenStudy (whpalmer4):

If I might make a suggestion, how about \( u = \ln(x^2)\)

OpenStudy (jaredstone4):

would it work if I changed the function to 2ln(x) /x , then pulled the 2 in front of the integral, and used u = ln(x)?

OpenStudy (mathmale):

\[2\int\limits_{?}^{?} \frac{ \ln x dx }{ x }\]Yes; that's it. ln x^2 = 2 ln x. You could move that coefficient 2 outside of the integral, ending up with

OpenStudy (mathmale):

Now try a u substitution.

OpenStudy (jaredstone4):

I put u = lnx, du/dx = 1/x, du = dx/x.

OpenStudy (jaredstone4):

And from there I know how to integrate it so I ended up with 4 as a final answer. Is that correct?

OpenStudy (mathmale):

I haven't tried obtaining the definite integral, but you sound so competent that I'm willing to take your word for it.

OpenStudy (mathmale):

Note that this is not the only way in which you could integrate. whpalmer suggested using u=ln (x^2), and upon thinking about that, I believe integration by parts would work here.

OpenStudy (mathmale):

But the method you and I used is probably the simplest.

OpenStudy (jaredstone4):

Haha well thank you, I appreciate that. And yes, I think it would work, too. This question is on a review packet from previous chapters and integration by parts I actually just learned this week so I figured it'd be best to use previous techniques.

OpenStudy (mathmale):

It gets to be a game after a while, a fun game, to figure out which methods work best in which situations. Do let me know if you have further qu. later.

OpenStudy (jaredstone4):

Fun is relative lol. I do have another one right now if you're willing. Which of the following are antiderivatives of f(x) = \[\frac{ 3\ln^2x }{ 2x}\]? I. \[\frac{ \ln^3x }{ 2 }\] II. \[\frac{ 1 }{ x } lnx^2 - \ln ^2x\] III. \[\frac{ 11 + \ln^3x }{ 2 }\]

OpenStudy (whpalmer4):

If you go with my suggestion, it's \[du = \frac{2}{x} \,dx\]\[\frac{1}{2}\int u \,du = \frac{1}{4}u^2+C\] then undo the substitution and evaluate.

OpenStudy (jaredstone4):

Answer choices: A) I only, B) II only, C) I and II only, D) I and III only, E) II and III only Thanks for the suggestion, if I had more time to experiment with IBP I'd try it :)

OpenStudy (mathmale):

Jared: Would you mind posting this new problem separately?

OpenStudy (jaredstone4):

Sure

OpenStudy (mathmale):

Just type the function to be integrated, not the possible answer choices.